# Edward Witten

### v1

@@ 8.4 Rényi Entropy and Generalized Entropy of a Black Hole
We see that in order $G^{0}$, the correction to the area contributes only the finite term $b_{1}/4$ to the entropy. But in order $G$, there is a divergent term $-f^{\prime}\Lambda^{2}b_{1}$. Therefore, in two-loop order and higher, the correction to the horizon area plays a role in the cancellation of divergences that makes the generalized entropy finite.
@@ 9.1 Heuristic Motivation
With this aim, we go back to the original spacetime of fig. 21(a), before any manipulations by the two observers. The extended Schwarzschild solution has a time translation symmetry and an $SO(3)$ symmetry that rotates the polar angles (which are not shown in the Penrose diagram). The Penrose diagram exhibits an additional ${\mathbb{Z}}_{2}\times{\mathbb{Z}}_{2}$ symmetry, where one ${\mathbb{Z}}_{2}$ is a spatial reflection that exchanges the left and right ends of the figure, and the other ${\mathbb{Z}}_{2}$ is a time-reversal symmetry that exchanges the past and future. The bifurcation surface $\Sigma$ is invariant under ${\mathbb{Z}}_{2}\times{\mathbb{Z}}_{2}$, and this ensures that $\Sigma$ is an extremal surface, meaning a surface whose area is invariant to first order if it is slightly displaced in any direction. This extremal property can also be deduced from the fact that $\Sigma$ is the fixed point set of the time-translation symmetry.
@@
The bifurcation surface $\Sigma$ is at $U=V=0$, and from the formula (253), is is evident that $\left.\frac{\partial r}{\partial U}\right|_{U=V=0}=\left.\frac{\partial r}{\partial V}\right|_{U=V=0}=0$. Since the area is $4\pi r^{2}$, the fact that $r$ is stationary at the horizon again shows that $\Sigma$ is an extremal surface.
@@ 9.2 The Original Setting
The RT formula RT was originally formulated as a statement about the entropy of a region in the boundary conformal field theory. We recall that according to this duality, a particular $d$-dimensional CFT, formulated on a $d$-manifold $Y$, has the same content as a $D=d+1$-dimensional gravitational theory484848 In this article, we have generally discussed black holes in the familiar $D=4$ world. Extension to other values of $D$ would generally not have added much. However, in analyzing the RT formula, we will consider arbitrary $D$ for several reasons. One reason is that in string theory, there are important examples with various values of $D$. Another reason is that the case $D=3$, $d=2$ is illuminating and easy to visualize. Yet another reason is that if one wants to use the RT formula to learn about gravity, one might consider $D=4$, $d=3$ to be the most natural case, but if one wants to use it to study entanglement entropy in quantum field theory, then one might be particularly interested in $D=5$, $d=4$. The cases $D=4,d=3$ and $D=3,d=2$ are also important in some condensed matter applications. So in short it seems artificial to limit a discussion of the RT formula to a particular value of $D$. formulated on an asymptotically locally AdS manifold $X$ whose conformal boundary is $Y$. In principle, for given $Y$, one must take into account all possible choices of $X$, but in the sort of simple applications of the RT formula that we will discuss, there is a particular $X$ that dominates for small $G$. In fact (as in RT ), the main example that we will discuss here is that $Y$ is the $d$-dimensional Einstein static universe $S^{d-1}\times{\mathbb{R}}$, and $X$ is ${\rm AdS}_{D}$, Anti de Sitter space of dimension $D=d+1$. As in the original formulation of the RT formula, we will assume a time-reversal symmetry $t\to-t$ that leaves fixed the initial value surfaces ${\mathcal{S}}\subset X$ and ${\mathcal{S}}^{\prime}\subset Y$, respectively of dimension $d$ and $d-1$. Here ${\mathcal{S}}^{\prime}$ is the conformal boundary of ${\mathcal{S}}$. (To be more precise, we formulate the CFT on a manifold $Y$ with a time-reversal symmetry, and we assume – as is true in simple examples, including those we will discuss – that the relevant $X$ also possesses this symmetry.) Now, let ${\sf A}$ be some region in ${\mathcal{S}}^{\prime}$. We assume that the boundary of ${\sf A}$ is an embedded $d-2$-dimensional manifold $\partial{\sf A}\subset{\mathcal{S}}^{\prime}$. For some given state $\Psi$ of the quantum fields on $Y$, let $\rho_{\sf A}$ be the density matrix appropriate for measurements in the region ${\sf A}$. The original version of the RT formula answered the following question: What is the von Neumann entropy $S_{\rm vN}(\rho_{\sf A})$? The proposed answer was that this entropy is given by a formula analogous to the Bekenstein-Hawking entropy formula
@@
where $\Sigma_{\sf A}$ is a particular surface known as the RT surface of ${\sf A}$, and the omitted terms are ${\mathcal{O}}(1)$ for small $G$. The definition of $\Sigma_{\sf A}$ is as follows: among all $d-1$-dimensional surfaces in ${\mathcal{S}}$ whose boundary coincides with $\partial A$ and which are also homologous to ${\sf A}$, $\Sigma_{\sf A}$ is the one whose area $A(\Sigma_{\sf A})$ is smallest. To be more precise, this “area” is really a renormalized area, as discussed at the end of this section, and as $\Sigma_{\sf A}$ has dimension $d-1$, it is really an area in the $d-1$-dimensional sense (thus it is only an ordinary area if $D=4$, $d=3$). The statement that $\Sigma_{\sf A}$ is homologous to ${\sf A}$ means that $\Sigma_{\sf A}$ and ${\sf A}$ together are the boundary of a region in ${\mathcal{S}}$. There is no explicit dependence on the quantum state $\Psi$ in the formula (256), so if true this formula implies that for small $G$, the leading contribution to the entropy depends on $\Psi$ only to the extent that the geometry depends on $\Psi$. When it is clear what density matrix is intended, we abbreviate $S_{\rm vN}(\rho_{\sf A})$ as $S_{\sf A}$.
@@
In this situation, $\Psi_{\rm HHI}$ is a pure state of the combined system ${\sf A}{\sf B}{\sf C}$. Hence $S_{\sf A}={\mathcal{S}}_{{\sf B}{\sf C}}$, and subadditivity of entropy $S_{\sf B}+S_{\sf C}\geq S_{{\sf B}{\sf C}}$ tells us that
@@
We have here considered only the leading divergence in the entanglement entropy. But in fact the parallelism between the predictions of field theory on the boundary and the RT formula in the bulk continues for subleading divergences, which in general can be matched between the boundary and bulk descriptions. For $d=2$, there is only the logarithmic divergence in the entanglement entropy that we have analyzed, but for $d>2$, in general the expansion of $S_{\sf A}$ for small $\varepsilon$ has a succession of less singular terms proportional to lower powers of $\frac{1}{\varepsilon}$, with a $\log\frac{1}{\varepsilon}$ term at the end for even $d$. (In some cases, depending on the dimensions of relevant operators in a particular theory, there can be a logarithm also for odd $d$.) When there is no logarithmic term, it is possible, by subtracting the power law divergences, to define a Weyl-invariant regularized version of the entanglement entropy. When there is a logarithm, the finite part of the entanglement entropy has a Weyl anomaly, as explained for $d=2$ at the end of section 8.2. Divergences and conformal anomalies in areas or volumes of extremal surfaces asymptotic to the boundary of an asymptotically locally AdS manifold were originally studied with a different motivation KS ; GW .
@@ 9.3 Derivation
Now, in AdS/CFT duality, consider the gravitational path integral in which we sum over bulk manifolds that have ${\mathcal{S}}^{\prime}\times S^{1}_{\beta}$ as their conformal boundary, If there is a unique manifold $X$ that dominates this sum as $G\to 0$, the time-reversal symmetry of the conformal boundary must extend to a time-reversal symmetry of $X$. Assuming that this is the case, [^50] The alternative is that rather than a single manifold that dominates for small $G$, there might be a pair of dominant manifolds exchanged by the time-reversal symmetry of the boundary conditions. In that case, the time-reversal symmetry of the density matrix is spontaneously broken as $G\to 0$. the time-reversal symmetry of $X$ will have a fixed point set ${\mathcal{S}}$ whose conformal boundary will be the fixed point set ${\mathcal{S}}^{\prime}\times(p\cup p^{\prime})$ of the boundary. Let $t_{\sf E}$ be a Euclidean time coordinate that vanishes along ${\mathcal{S}}$ and is odd under time-reversal. The time-reversal symmetry of the Euclidean manifold $X$ ensures that by defining $t=-{\rm i}t_{\sf E}$ and taking $t$ to be real, we get a real Lorentz signature spacetime $X_{L}$. (This analytic continuation leads to a real metric by the same reasoning as in section 6.7, though now we are proceeding in reverse, starting in Euclidean signature and continuing to Lorentz signature.) We can view $X_{L}$ as the Lorentz signature spacetime that evolves from the initial value surface ${\mathcal{S}}$.
@@
Let us consider the $n$-fold cover $X_{n}$ as a candidate for what might be the dominant manifold $X_{[n]}$. As $X_{n}$ is locally isomorphic to $X$ but covers it $n$ times, its classical action is $I(X_{n})=nI(X)$. Therefore, if $X_{n}$ is the dominant manifold in the computation of ${\rm Tr}\,\rho^{n}$, then eqn (271) tells us that $R_{n}(\rho)$ vanishes in order $1/G$. If so, then by analytic continuation, all Rényi and von Neumann entropies vanish. Could there be a more dominant contribution to ${\rm Tr}\,\rho^{n}$? If there is another manifold that is more dominant than $X_{n}$, its action must be less then $I(X_{n})=nI(X)$, and in that case, eqn. (271) tells us that $R_{n}(\rho)<0$ in order $1/G$. For sufficiently small $G$, this contradicts the positivity of Rényi entropies (eqn. (183)). So $X_{n}$ is the dominant manifold, and $R_{n}(\rho)$ vanishes in order $1/G$. By analytic continuation in $n$, this is true for all Rényi entropies $R_{\alpha}(\rho)$ and for the von Neumann entropy.
@@
If $\rho$ is a density matrix that is not purely thermal, for example if we deform away from a thermal density matrix as in eqn. (265), the metric of the dominant Euclidean manifold will change, but at least for a sufficiently small perturbation, the topology will be the same. Thus for a whole open set in the space of density matrices, we expect the dominant manifold $X$ to be topologically ${\mathcal{S}}^{\prime}\times D$. Moreover, for a time-reversal invariant density matrix that is sufficiently close to a thermal one, we expect the time-reversal symmetry to remain unbroken, so the dominant metric on $X=X_{[1]}$ will possess the time-reversal symmetry of the boundary. The fixed point set of this time-reversal symmetry will be topologically ${\mathcal{S}}={\mathcal{S}}^{\prime}\times{\mathbb{R}}$, as in the thermal case. Here ${\mathcal{S}}$ is the time-reversal invariant Cauchy hypersurface that was introduced earlier in constructing the Lorentz signature spacetime $X_{L}$ (but now we know the topology of ${\mathcal{S}}$ because the density matrix is close to a thermal one). The RT formula says that there should be a minimum area surface $\Sigma\subset{\mathcal{S}}$ – the RT surface – whose area determines the entropy of $\rho$ in leading order:
@@
The analog here is the following. As long as $n\not=1$, there is no meaningful condition to place on the locus $\Sigma_{[n]}$ of a conical singularity in $X^{\prime}_{[n]}$. The geometry of $X^{\prime}_{[n]}$ away from the conical singularity is determined by the Einstein equations. There is no way to move the singularity without changing the geometry and we are not free to change the geometry as it is determined by the Einstein equations. If one considers the conical singularity to be sourced by a cosmic brane, one would say that the Einstein equations including this source determine the spacetime geometry and imply that the energy-momentum tensor of the brane is conserved. In the limit that $n\to 1$, however, there is no conical singularity, the cosmic brane is not sourcing anything, and one does need to impose the condition that its energy-momentum tensor is conserved. Generalizing the geodesic equation for a test particle, the condition of conservation of the stress tensor for a brane in an ambient spacetime is that its worldvolume has extremal area. Analogously to the EIH result and its counterpart for the Moon, it is shown in LM that in the limit $n\to 1$, the Einstein equations for $X^{\prime}_{[n]}$ go over to the Einstein equations for $X$ together with the condition that $\Sigma$ is a surface of extremal area.
@@
${\rm Tr}\,\rho_{\sf A}^{n}=\exp(-I(X_{[n]}).$ (279)
(The normalizing factor $\exp(nI(X_{[1]})$ in eqn. (270) is, according to AdS/CFT duality, equal to $(|\Psi|^{2})^{n}$, so it equals 1 if $\Psi$ is normalized.) Now, as before, we assume that $X_{[n]}$ is invariant under the ${\mathbb{Z}}_{[n]}$ replica symmetry and we define $X^{\prime}_{[n]}=X_{[n]}/{\mathbb{Z}}_{n}$. $X^{\prime}_{[n]}$ will have a conical singularity, with cone angle $2\pi/n$, on a codimension two manifold $\Sigma_{[n]}$ whose conformal boundary is $\partial{\sf A}$. The conformal boundary of $\Sigma_{[n]}$ is $\partial{\sf A}$ because the conformal boundary of $X^{\prime}_{[n]}$ is $Y^{\prime}_{[n]}=Y_{[n]}/{\mathbb{Z}}_{n}$, and $Y^{\prime}_{[n]}$ has a conical singularity along $\partial{\sf A}$. As in the previous discussion, the formula (279) for ${\rm Tr}\,\rho_{\sf A}^{n}$ can be analytically continued in $n$ and leads to a simple answer in the limit $n\to 1$. Following the same logic as before, one finds a formula for the von Neumann entropy of $\rho_{\sf A}$ in terms of the (renormalized) area of a surface $\Sigma$ that is the limit of $\Sigma_{[n]}$ for $n\to 1$:
@@
In general, not much is known about constructing candidate manifolds $X_{\mu,n}$ for integer $n>1$ associated to candidate RT surfaces $\Sigma_{\mu}$. In a few cases, this has been done. For instance, in the example shown in fig, 25 with two candidate RT surfaces $\Sigma_{\mu}$, $\mu=1,2$, manifolds $X_{\mu,2}$ were explicitly described and studied in Headrick , and this has been extended to $X_{\mu,n}$ BDHM ..

### v4

@@ 8.4 Rényi Entropy and Generalized Entropy of a Black Hole
We see that in order $G^{0}$, the correction to the area contributes only the finite term $b_{1}/4$ to the entropy. But in order $G$, there is a divergent term $-f^{\prime}\Lambda^{2}b_{1}/4$. Therefore, in two-loop order and higher, the correction to the horizon area plays a role in the cancellation of divergences that makes the generalized entropy finite.
@@ 9.1 Heuristic Motivation
With this aim, we go back to the original spacetime of fig. 21(a), before any manipulations by the two observers. The extended Schwarzschild solution has a time translation symmetry and an $SO(3)$ symmetry that rotates the polar angles (which are not shown in the Penrose diagram). The Penrose diagram exhibits an additional ${\mathbb{Z}}_{2}\times{\mathbb{Z}}_{2}$ symmetry, where one ${\mathbb{Z}}_{2}$ is a spatial reflection that exchanges the left and right ends of the figure, and the other ${\mathbb{Z}}_{2}$ is a time-reversal symmetry that exchanges the past and future. The bifurcation surface $\Sigma$ is invariant under ${\mathbb{Z}}_{2}\times{\mathbb{Z}}_{2}$, and this ensures that $\Sigma$ is an extremal surface, meaning a surface whose area is invariant to first order if it is slightly displaced in any direction. This extremal property can also be deduced from the fact that $\Sigma$ is the fixed point set of the time-translation symmetry. [^45] Because time-translation symmetry rescales $U$ and $V$, it implies that the horizon area is independent of $U$ at $V=0$ and independent of $V$ at $U=0$. (We see this explicitly in eqn. (163): if $U$ or $V$ vanishes, then $r=2GM$.) Hence at $U=V=0$, the $U$ and $V$ derivatives of the horizon area both vanish, so the bifurcation surface at $U=V=0$ is an extremal surface.
@@
The bifurcation surface $\Sigma$ is at $U=V=0$, and from the formula (253), it is evident that $\left.\frac{\partial r}{\partial U}\right|_{U=V=0}=\left.\frac{\partial r}{\partial V}\right|_{U=V=0}=0$. Since the area is $4\pi r^{2}$, the fact that $r$ is stationary at the horizon again shows that $\Sigma$ is an extremal surface.
@@ 9.2 The Original Setting
The RT formula RT was originally formulated as a statement about the entropy of a region in the boundary conformal field theory. We recall that according to this duality, a particular $d$-dimensional CFT, formulated on a $d$-manifold $Y$, has the same content as a $D=d+1$-dimensional gravitational theory494949 In this article, we have generally discussed black holes in the familiar $D=4$ world. Extension to other values of $D$ would generally not have added much. However, in analyzing the RT formula, we will consider arbitrary $D$ for several reasons. One reason is that in string theory, there are important examples with various values of $D$. Another reason is that the case $D=3$, $d=2$ is illuminating and easy to visualize. Yet another reason is that if one wants to use the RT formula to learn about gravity, one might consider $D=4$, $d=3$ to be the most natural case, but if one wants to use it to study entanglement entropy in quantum field theory, then one might be particularly interested in $D=5$, $d=4$. The cases $D=4,d=3$ and $D=3,d=2$ are also important in some condensed matter applications. So in short it seems artificial to limit a discussion of the RT formula to a particular value of $D$. formulated on an asymptotically locally AdS manifold $X$ whose conformal boundary is $Y$. In principle, for given $Y$, one must take into account all possible choices of $X$, but in the sort of simple applications of the RT formula that we will discuss, there is a particular $X$ that dominates for small $G$. In fact (as in RT ), the main example that we will discuss here is that $Y$ is the $d$-dimensional Einstein static universe $S^{d-1}\times{\mathbb{R}}$, and $X$ is ${\rm AdS}_{D}$, Anti de Sitter space of dimension $D=d+1$. As in the original formulation of the RT formula, we will assume a time-reversal symmetry $t\to-t$ that leaves fixed the initial value surfaces ${\mathcal{S}}\subset X$ and ${\mathcal{S}}^{\prime}\subset Y$, respectively of dimension $d$ and $d-1$. Here ${\mathcal{S}}^{\prime}$ is the conformal boundary of ${\mathcal{S}}$. (To be more precise, we formulate the CFT on a manifold $Y$ with a time-reversal symmetry, and we assume – as is true in simple examples, including those we will discuss – that the relevant $X$ also possesses this symmetry.) Now, let ${\sf A}$ be some region in ${\mathcal{S}}^{\prime}$. We assume that the boundary of ${\sf A}$ is an embedded $d-2$-dimensional manifold $\partial{\sf A}\subset{\mathcal{S}}^{\prime}$. For some given state $\Psi$ of the quantum fields on $Y$, let $\rho_{\sf A}$ be the density matrix appropriate for measurements in the region ${\sf A}$. The original version of the RT formula answered the following question: What is the von Neumann entropy $S_{\rm vN}(\rho_{\sf A})$? The proposed answer was that this entropy is given by a formula analogous to the Bekenstein-Hawking entropy formula
@@
where $\Sigma_{\sf A}$ is a particular surface known as the RT surface of ${\sf A}$, and the omitted terms are ${\mathcal{O}}(1)$ for small $G$. The definition of $\Sigma_{\sf A}$ is as follows: among all $d-1$-dimensional surfaces in ${\mathcal{S}}$ whose boundary coincides with $\partial{\sf A}$ and which are also homologous to ${\sf A}$, $\Sigma_{\sf A}$ is the one whose area $A(\Sigma_{\sf A})$ is smallest. To be more precise, this “area” is really a renormalized area, as discussed at the end of this section, and as $\Sigma_{\sf A}$ has dimension $d-1$, it is really an area in the $d-1$-dimensional sense (thus it is only an ordinary area if $D=4$, $d=3$). The statement that $\Sigma_{\sf A}$ is homologous to ${\sf A}$ means that $\Sigma_{\sf A}$ and ${\sf A}$ together are the boundary of a region in ${\mathcal{S}}$. There is no explicit dependence on the quantum state $\Psi$ in the formula (256), so if true this formula implies that for small $G$, the leading contribution to the entropy depends on $\Psi$ only to the extent that the geometry depends on $\Psi$. When it is clear what density matrix is intended, we abbreviate $S_{\rm vN}(\rho_{\sf A})$ as $S_{\sf A}$.
@@
In this situation, $\Psi_{\rm HHI}$ is a pure state of the combined system ${\sf A}{\sf B}{\sf C}$. Hence $S_{\sf A}=S_{{\sf B}{\sf C}}$, and subadditivity of entropy $S_{\sf B}+S_{\sf C}\geq S_{{\sf B}{\sf C}}$ tells us that
@@
We have here considered only the leading divergence in the entanglement entropy. But in fact the parallelism between the predictions of field theory on the boundary and the RT formula in the bulk continues for subleading divergences, which in general can be matched between the boundary and bulk descriptions. For $d=2$, the entanglement entropy has only the logarithmic divergence that we have analyzed, but for $d>2$, in general the expansion of $S_{\sf A}$ for small $\varepsilon$ has a succession of less singular terms proportional to lower powers of $\frac{1}{\varepsilon}$, with a $\log\frac{1}{\varepsilon}$ term at the end for even $d$. (In some cases, depending on the dimensions of relevant operators in a particular theory, there can be a logarithm also for odd $d$.) When there is no logarithmic term, it is possible, by subtracting the power law divergences, to define a Weyl-invariant regularized version of the entanglement entropy. When there is a logarithm, the finite part of the entanglement entropy has a Weyl anomaly, as explained for $d=2$ at the end of section 8.2. Divergences and conformal anomalies in areas or volumes of extremal submanifolds asymptotic to the boundary of an asymptotically locally AdS manifold were originally studied with a different motivation KS ; GW .
@@ 9.3 Derivation
Now, in AdS/CFT duality, consider the gravitational path integral in which we sum over bulk manifolds that have ${\mathcal{S}}^{\prime}\times S^{1}_{\beta}$ as their conformal boundary. If there is a unique manifold $X$ that dominates this sum as $G\to 0$, the time-reversal symmetry of the conformal boundary must extend to a time-reversal symmetry of $X$. Assuming that this is the case, [^51] The alternative is that rather than a single manifold that dominates for small $G$, there might be a pair of dominant manifolds exchanged by the time-reversal symmetry of the boundary conditions. In that case, the time-reversal symmetry of the density matrix is spontaneously broken as $G\to 0$. the time-reversal symmetry of $X$ will have a fixed point set ${\mathcal{S}}$ whose conformal boundary will be the fixed point set ${\mathcal{S}}^{\prime}\times(p\cup p^{\prime})$ of the boundary. Let $t_{\sf E}$ be a Euclidean time coordinate that vanishes along ${\mathcal{S}}$ and is odd under time-reversal. The time-reversal symmetry of the Euclidean manifold $X$ ensures that by defining $t=-{\rm i}t_{\sf E}$ and taking $t$ to be real, we get a real Lorentz signature spacetime $X_{L}$. (This analytic continuation leads to a real metric by the same reasoning as in section 6.7, though now we are proceeding in reverse, starting in Euclidean signature and continuing to Lorentz signature.) We can view $X_{L}$ as the Lorentz signature spacetime that evolves from the initial value surface ${\mathcal{S}}$.
@@
Let us consider the $n$-fold cover $X_{n}$ as a candidate for what might be the dominant manifold $X_{[n]}$. As $X_{n}$ is locally isomorphic to $X$ but covers it $n$ times, its classical action is $I(X_{n})=nI(X)$. Therefore, if $X_{n}$ is the dominant manifold in the computation of ${\rm Tr}\,\rho^{n}$, then eqn (271) tells us that $R_{n}(\rho)$ vanishes in order $1/G$. If so, then by analytic continuation, all Rényi and von Neumann entropies vanish. Could there be a more dominant contribution to ${\rm Tr}\,\rho^{n}$? If there is another manifold that is more dominant than $X_{n}$, its action must be less then $I(X_{n})=nI(X)$, and in that case, eqn. (271) tells us that $R_{n}(\rho)<0$ in order $1/G$. For sufficiently small $G$, this contradicts the positivity of Rényi entropies (eqn. (183)). So $X_{n}$ is the dominant manifold, and $R_{n}(\rho)$ vanishes in order $1/G$. By analytic continuation in $n$, this is true for all Rényi entropies $R_{\alpha}(\rho)$ and for the von Neumann entropy. [^53] More explicitly, since $R_{n}(\rho)$ is non-increasing as a function of $n$ (eqn. (184)) and non-negative, its vanishing for any given $n$ implies vanishing for all $n^{\prime}<n$.
@@
If $\rho$ is a density matrix that is not purely thermal, for example if we define $\rho$ by deforming away from a thermal density matrix as in eqn. (265), the metric of the dominant Euclidean manifold will change, but at least for a sufficiently small perturbation, the topology will be the same. Thus for a whole open set in the space of density matrices, we expect the dominant manifold $X$ to be topologically ${\mathcal{S}}^{\prime}\times D$. Moreover, for a time-reversal invariant density matrix that is sufficiently close to a thermal one, we expect the time-reversal symmetry to remain unbroken, so the dominant metric on $X=X_{[1]}$ will possess the time-reversal symmetry of the boundary. The fixed point set of this time-reversal symmetry will be topologically ${\mathcal{S}}={\mathcal{S}}^{\prime}\times{\mathbb{R}}$, as in the thermal case. Here ${\mathcal{S}}$ is the time-reversal invariant Cauchy hypersurface that was introduced earlier in constructing the Lorentz signature spacetime $X_{L}$ (but now we know the topology of ${\mathcal{S}}$ because the density matrix is close to a thermal one). The RT formula says that there should be a minimum area surface $\Sigma\subset{\mathcal{S}}$ – the RT surface – whose area determines the entropy of $\rho$ in leading order:
@@
The analog here is the following. As long as $n\not=1$, there is no meaningful condition on the position of the locus $\Sigma_{[n]}$ of a conical singularity in $X^{\prime}_{[n]}$. The geometry of $X^{\prime}_{[n]}$ away from the conical singularity is determined by the Einstein equations. There is no way to move the singularity without changing the geometry and we are not free to change the geometry as it is determined by the Einstein equations. If one considers the conical singularity to be sourced by a cosmic brane, one would say that the Einstein equations including this source determine the spacetime geometry and imply that the energy-momentum tensor of the brane is conserved. In the limit that $n\to 1$, however, there is no conical singularity, the cosmic brane is not sourcing anything, and one does need to impose the condition that its energy-momentum tensor is conserved. Generalizing the geodesic equation for a test particle, the condition of conservation of the stress tensor for a brane in an ambient spacetime is that its worldvolume has extremal area. Analogously to the EIH result and its counterpart for the Moon, it is shown in LM that in the limit $n\to 1$, the Einstein equations for $X^{\prime}_{[n]}$ go over to the Einstein equations for $X$ together with the condition that $\Sigma$ is a surface of extremal area.
@@
${\rm Tr}\,\rho_{\sf A}^{n}=\exp(-I(X_{[n]})).$ (279)
(The normalizing factor $\exp(nI(X_{[1]})$ in eqn. (270) is, according to AdS/CFT duality, equal to $(|\Psi|^{2})^{-n}$, so it equals 1 if $\Psi$ is normalized.) Now, as before, we assume that $X_{[n]}$ is invariant under the ${\mathbb{Z}}_{[n]}$ replica symmetry and we define $X^{\prime}_{[n]}=X_{[n]}/{\mathbb{Z}}_{n}$. $X^{\prime}_{[n]}$ will have a conical singularity, with cone angle $2\pi/n$, on a codimension two manifold $\Sigma_{[n]}$ whose conformal boundary is $\partial{\sf A}$. The conformal boundary of $\Sigma_{[n]}$ is $\partial{\sf A}$ because the conformal boundary of $X^{\prime}_{[n]}$ is $Y^{\prime}_{[n]}=Y_{[n]}/{\mathbb{Z}}_{n}$, and $Y^{\prime}_{[n]}$ has a conical singularity along $\partial{\sf A}$. As in the previous discussion, the formula (279) for ${\rm Tr}\,\rho_{\sf A}^{n}$ can be analytically continued in $n$ and leads to a simple answer in the limit $n\to 1$. Following the same logic as before, one finds a formula for the von Neumann entropy of $\rho_{\sf A}$ in terms of the (renormalized) area of a surface $\Sigma$ that is the limit of $\Sigma_{[n]}$ for $n\to 1$:
@@
In general, not much is known about constructing candidate manifolds $X_{\mu,n}$ for integer $n>1$ associated to candidate RT surfaces $\Sigma_{\mu}$. In a few cases, this has been done. For instance, in the example shown in fig, 25 with two candidate RT surfaces $\Sigma_{\mu}$, $\mu=1,2$, manifolds $X_{\mu,2}$ were explicitly described and studied in Headrick , and this has been extended to $X_{\mu,n}$ BDHM .